Zodwa P=50-2Q Lihle P=60-3Q

Deduce the market quantity identity for ice cream


1
Expert's answer
2022-03-27T11:04:26-0400

"P=50-2Q,"

"P=60-3Q,"

"50-2Q=60-3Q,"

"Q=10,"

"P=30."